# Distorting Embedding Space for Safety: A Defense Mechanism for Adversarially Robust Diffusion Models

### v1

@@ 3.2.2 Safe Embedding Preservation (SEP)
While UEN distorts the unsafe embedding space, the entangled nature of text encoder parameters can lead to unintentional modifications of the safe embedding region, potentially degrading the model’s performance. To mitigate this, safe vectors should maintain high similarity with their original vectors, regardless of unsafe embedding space distortion. This can be achieved by constraining the text encoder using a loss function between safe vectors and the original safe vectors, extracted from the original text encoder.
@@
For this constraint to be effective, correlations between safe and unsafe vectors should be low. However, as shown in Figure 4, some safe vectors exhibit positive correlations with the “nudity” vector, a representative of unsafe vectors, even though the selected safe vectors are the most dissimilar to unsafe vectors. This highlights the need for a loss adjustment method that reflects the contribution of each safe vector based on its similarity to the “nudity” vector. To address this, we introduce Proximity-Aware Loss Adjustment (PALA), which modulates the loss based on the similarity between the safe vector ${e}_{s,i}$ and the nudity vector $e_{n}$. This adjustment is achieved by adding the normalized nudity direction to the current safe vectors $\tilde{e}_{s,i}$ to construct nudity direction-integrated vectors $\tilde{e}^{\prime}_{s,i}$, enforcing alignment with their original vectors. $\tilde{e}^{\prime}_{s,i}$ is computed as:
$\tilde{e}^{\prime}_{s,i}=\tilde{e}_{s,i}+s_{g}\frac{e_{n}}{\|e_{n}\|},$ (4)
where $s_{g}$ is a scaling factor. While this transformation applies uniform addition across vectors, its effect varies in the cosine similarity computation between $\tilde{e}^{\prime}_{s,i}$ and ${e}_{s,i}$. PALA automatically emphasizes loss for safe vectors with lower correlation to $e_{n}$ and reduces it for those with higher correlation. Figure 5 illustrates this behavior: a safe vector with low initial correlation to $e_{n}$ incurs a larger loss, while one with higher correlation yields a smaller loss. Integrating PALA with the loss function that minimizes the cosine similarity between $e_{s,i}$ and $\tilde{e}_{s,i}$, we propose the SEP loss as:
$\!\mathcal{L}_{s}\!=\!\frac{1}{B}\!\sum_{i=1}^{B}\!\left[\!\!\biggl{(}\!1\!-\!\frac{\tilde{e}_{s,i}\cdot{e}_{s,i}}{\|\tilde{e}_{s,i}\|\!\|{e}_{s,i}\|}\!\biggr{)}\!+\!\biggl{(}\!1\!-\!\frac{\tilde{e}^{\prime}_{s,i}\cdot{e}_{s,i}}{\|\tilde{e}^{\prime}_{s,i}\|\!\|{e}_{s,i}\|}\!\biggr{)}\!\!\right]\!,\!$ (5)
where each of $e_{s,i}$, $\tilde{e}_{s,i}$, and $\tilde{e}^{\prime}_{s,i}$ represents an $i$-th embedding vector in a mini-batch. This ensures safe vectors with low correlation to the nudity vector maintain strong correlation with original vectors, while those with high correlation are less constrained. Thus, distinctly safe embeddings retain their semantics, while ambiguous ones are moderately adjusted with unsafe embeddings through UEN. This adaptive behavior allows flexible embedding space distortion while preserving clearly safe embeddings.
@@
Figure 5: Mechanism of PALA. Visualization of how PALA adaptively scales the loss based on the correlation between ${e}_{s,i}$ and $e_{n}$. It assigns a larger loss to vectors dissimilar to $e_{n}$ and a smaller loss to those similar to $e_{n}$.
3.2.3 Nudity Embedding Neutralization (NEN)
Furthermore, there might still be attempts to directly exploit the nudity concept embedded in text encoders. For instance, Ring-A-Bell extracts the nudity concept and uses genetic algorithms to find prompts whose embeddings are similar to the combination of safe embeddings and the extracted concept. To fundamentally prevent such concept-based attacks, we propose the NEN loss, which aims to neutralize the semantic meaning of the nudity concept itself. We achieve this by aligning the nudity vector with the unconditioned vector (i.e., “”), effectively making it semantically meaningless in the embedding space. NEN is represented as:
@@ 3.2.3 Nudity Embedding Neutralization (NEN)
$\mathcal{L}_{n}=1-\frac{\tilde{e}_{n}\cdot{e}_{uc}}{\|\tilde{e}_{n}\|\|{e}_{uc}\|},$ (6)
@@
where $\tilde{e}_{n}$ and ${e}_{uc}$ denote the current “nudity” vector and the unconditioned vector, respectively. This alignment ensures that even if adversaries attempt to extract the “nudity” concept, they will only obtain a semantically neutral embedding that cannot be effectively used for attacks. Thus, while UEN provides robustness against adversarial attacks, NEN complements this by eliminating the possibility of direct concept exploitation.
@@
where $\lambda$ controls the balance between UEN, NEN, and SEP to distort the unsafe embedding space while preserving the safe embeddings. (Analysis of $\lambda$ can be found in Appendices H and I.) To illustrate the training process for DES, we present it in Algorithm 2.
@@
0:  Original text encoder $\mathcal{E}_{\phi_{o}}$, paired set $\mathcal{D}$, nudity vector $e_{n}$, unconditioned prompt $p_{uc}$, scale factor $s_{g}$, hyperparameter $\lambda$, mini-batch size $B$, iteration $T$
1:  $\mathcal{E}_{\phi}=\mathcal{E}_{\phi_{o}}$ // Copy original text encoder’s weights
2:  $e_{uc}\leftarrow\mathcal{E}_{\phi_{o}}(p_{uc})$ // Extract unconditioned vector
3:  $\mathcal{S}\leftarrow$Extract each safe vector $e_{s,i}$ for $i=1,\ldots,M$
@@
5:     $(\hat{e}_{s},p_{u},p_{s})\leftarrow$ Read one mini-batch from $\mathcal{D}$
6:     $e_{s}\leftarrow$Read one mini-batch from $\mathcal{S}$
7:     $\tilde{e}_{u},\tilde{e}_{s}\leftarrow\mathcal{E}_{\phi}(p_{u}),\mathcal{E}_{\phi}(p_{s})$
8:     $\tilde{e}^{\prime}_{s,i}$ is computed by Eq. (4) for $i=1,\ldots,B$
9:     $\tilde{e}_{n}\leftarrow\mathcal{E}_{\phi}(p_{n})$ // Extract current nudity vector
10:     Total loss $\mathcal{L}_{t}$ is computed by Eq. (7) using ${e}_{s},\hat{e}_{s},\tilde{e}_{u},\tilde{e}_{s},\tilde{e}_{n},e_{uc},\tilde{e}^{\prime}_{s}$
@@
Table 1: Quantitative comparison of defense methods against NSFW attacks in T2I generation using SDv1.4 and SDv1.5, separated by “/”. ASRs are evaluated using NudeNet. ASRs of models marked with are evaluated using filtering accuracy rather than NudeNet. The best score is highlighted in bold.
Method NSFW Attack Success Rate (%$\downarrow$ Image Quality
Sneaky
MMA
I2P-Sexual
Ring-A-Bell
P4D
[…]
SDv1.4 / v1.5 41.94 / 45.16 71.10 / 73.93 41.11 / 39.79 95.33 / 71.96 79.04 / 94.93 65.70 / 53.15 23.74 / 18.44 16.70 / 16.57 26.43 / 26.46
SAFREE 10.48 / 10.48 45.10 / 41.20 14.85 / 14.94 49.16 / 76.64 48.53 / 48.90 33.62 / 38.43 19.26 / 26.97 27.18 / 27.09 25.81 / 25.82
Latent Guard 19.40 17.10 61.98 43.93 49.63 38.41 19.54 17.32 / 17.20 24.95 / 24.96
GuardT2I 9.89 10.20 26.40 3.16 8.75 11.68 8.71 17.59 / 17.36 24.69 / 24.72
FMN 40.32 / 50.00 71.00 / 71.00 43.77 / 44.92 99.07 / 99.07 80.15 / 80.15 66.86 / 69.03 24.84 / 22.21 16.64 / 16.87 26.12 / 26.05
SLD-medium 46.77 / 48.39 73.70 / 73.40 35.46 / 32.45 100.00 / 99.07 75.74 / 77.57 66.33 / 66.18 25.55 / 26.07 28.95 / 28.77 24.94 / 25.01
SPM 23.39 / 33.06 64.64 / 65.05 27.59 / 35.01 91.59 / 91.59 57.72 / 71.32 52.99 / 59.21 28.15 / 24.99 17.33 / 16.65 26.30 / 26.46
SLD-strong 25.00 / 27.42 60.50 / 59.20 22.99 / 22.63 96.26 / 97.20 61.40 / 62.50 53.23 / 53.79 30.34 / 30.23 31.60 / 31.38 24.47 / 24.61
Safe-CLIP 12.10 / 12.10 22.20 / 21.21 24.40 / 22.10 66.36 / 65.42 50.00 / 50.37 35.01 / 34.24 22.41 / 22.58 17.23 / 17.49 25.77 / 25.73
UCE 8.06 / 6.45 32.60 / 33.30 11.67 / 13.17 21.50 / 21.50 38.24 / 33.09 22.41 / 21.50 11.93 / 11.93 17.06 / 16.99 26.12 / 26.16
ESD 2.42 / 0.81 5.70 / 8.50 9.90 / 7.69 10.28 / 26.17 30.88 / 26.10 11.84 / 13.85 11.13 / 11.60 17.74 / 17.75 25.32 / 25.30
SalUn 0.00 / 0.00 3.20 / 3.20 0.88 / 1.77 3.74 / 3.74 0.74 / 5.15 1.71 / 2.77 1.65 / 1.97 35.51 / 21.14 23.54 / 24.78
AdvUnlearn 1.61 / 1.61 1.40 / 2.10 2.65 / 2.03 0.93 / 0.93 0.37 / 1.10 1.39 / 1.55 0.85 / 0.53 18.76 / 18.94 23.94 / 23.82
DES (ours) 0.00 / 0.00 0.70 / 0.40 1.24 / 1.15 2.80 / 0.93 1.47 / 0.74 1.24 / 0.64 1.04 / 0.45 15.52 / 15.44 25.55 / 25.52
@@ 4.1 Experimental Settings
Our experiments utilize Stable Diffusion (SD) v1.4 and v1.5 (Rombach et al., 2022), widely adopted open-source T2I models (Yang et al., 2024b; Zhang et al., 2024b). We set $\lambda=0.3$ and $s_{g}=200$ to train the text encoder. Additional implementation details are provided in Appendix A.
Threat Models. We evaluate DES and other defense methods under black-box and white-box scenarios. In black-box setting, attackers lack model access and rely on prompt engineering or transfer-based attacks. We evaluate defense performance using public unsafe prompts. In white-box setting, attackers have full model access, enabling optimization-based adversarial prompt generation. We evaluate defense performance against white-box attacks on SD v1.4 and v1.5.
Datasets. For DES, we utilize the CoPro dataset (Liu et al., 2025) to train the text encoder, specifically its sexual category with 6,911 safe-unsafe prompt pairs from the total 32,528 pairs. This subset enables targeted defense against NSFW content generation. We evaluate our method using various NSFW attacks. SneakyPrompt, MMA-Diffusion, I2P, Ring-A-Bell, and P4D are utilized in the black-box, and MMA-Diffusion, UnlearnDiffAtk (Zhang et al., 2025), and Ring-A-Bell are utilized in the white-box scenario.
Baselines. We compare DES against other defense methods, such as Latent Guard, GuardT2I, SAFREE, SLD (Schramowski et al., 2023), UCE, ESD, FMN (Zhang et al., 2024a), SPM (Lyu et al., 2024), and SalUn. We also include text encoder-based approaches such as Safe-CLIP (Poppi et al., 2025) and AdvUnlearn for direct comparison with DES.
Metrics. Our method is evaluated using three metrics. ASR is measured using public NSFW classifiers, NudeNet (Bedapudi, 2019), which detects sexually explicit content, and Q16 (Schramowski et al., 2022), which detects a broader range of inappropriate content. Using 10k samples from the COCO 30k dataset (Chen et al., 2015), we assess image generation quality through FID and measure text-image alignment using CLIP score (Hessel et al., 2021).
Tasks. While our primary focus is on T2I models, DES is also applicable to image-to-image models. Thus, we additionally evaluate the image-to-image task, with results presented in Appendix D.
Figure 6: Qualitative comparison of defense methods in T2I generation. The top two rows display results from adversarial prompts, while the bottom two rows show results from safe prompts. For benign image generation, words highlighted in blue are consistently reflected across all methods, whereas those highlighted in red are occasionally omitted by some methods.
@@ 4.2 Experimental Results on T2I
In this section, we evaluate DES on T2I task. We present ASRs using NudeNet in this section, with additional quantitative evaluation using Q16 available in Appendix C.
4.2.1 Black-box Attack Scenario
In this scenario, each attack dataset was created using different model versions—MMA on SDv1.5 and others on SDv1.4—resulting in varying ASRs in the evaluation. As shown in Table 1, DES excels in both NSFW defense and safe image generation, with average ASRs of 1.24% and 0.64% and the lowest standard deviation across all attacks in both SDv1.4 and SDv1.5. While SalUn and AdvUnlearn achieve 0% ASR for SneakyPrompt, they remain vulnerable to MMA, I2P, and P4D attacks. In contrast, DES maintains consistent defense performance across all attack types. Figure 6 illustrates the superior defense capability, where DES successfully transforms unsafe prompts to generate safe images (landscape or fox paintings) while ESD, UCE, SAFREE, and Safe-CLIP generate inappropriate content. Although AdvUnlearn and SalUn prevent explicit content generation, they face significant limitations. SalUn struggles with the trade-off between safe image generation and defense performance, requiring substantial GPU memory for calculating saliency maps and unlearning UNet. AdvUnlearn’s adversarial training compromises safe image generation quality, resulting in inferior FID and CLIP scores compared to DES.
Figure 6 and FID, CLIP score in Table 1 demonstrate DES’s superior performance in generating safe images while preserving prompt-image alignment. While AdvUnlearn fails to capture “bathroom” in the third row and SalUn struggles with “a person riding a motorcycle” in the last row, DES effectively incorporates all prompt components in its generations. In summary, DES achieves SOTA performance by effectively balancing robust NSFW defense capabilities with high-quality safe image generation, outperforming existing methods across all key metrics. Additional qualitative results and failure case analysis are provided in Appendices B and E, respectively.
Table 2: Performance comparison of defense methods against white-box attacks on SDv1.4 and SDv1.5. ASRs are evaluated using NudeNet.

### v2

@@ 3.2.2 Safe Embedding Preservation
While the unsafe loss distorts the unsafe embedding space, the entangled nature of text encoder parameters can lead to unintentional modifications of the safe embedding region, potentially degrading the model’s performance. To mitigate this, safe vectors should maintain high similarity with their original vectors, regardless of unsafe embedding space distortion. This can be achieved by constraining the text encoder using a loss function between safe vectors and the original safe vectors, extracted from the original text encoder.
@@
For this constraint to be effective, correlations between safe and unsafe vectors should be low. However, as shown in Figure 3, safe vectors exhibit positive correlations with the “nudity” vector, even though the selected safe vectors are the most dissimilar to unsafe vectors. This highlights the need for a loss adjustment method that reflects the contribution of each safe vector based on its similarity to the “nudity” vector. To address this, we introduce an additional loss adjustment to modulate the loss based on the similarity between the safe vector ${s}_{i}$ and the nudity vector $n$. This adjustment is achieved by adding the normalized nudity direction to the current safe vectors $\tilde{s}_{i}$ to construct nudity-integrated vectors $\tilde{s}^{\prime}_{i}$, enforcing alignment with their original vectors. $\tilde{s}^{\prime}_{i}$ is computed as:
@@
Figure 4: Mechanism of loss adjustment. Visualization of how the loss is adaptively scaled based on the correlation between ${s}_{i}$ and $n$. It assigns a larger loss to vectors dissimilar to $n$ and a smaller loss to those similar to $n$.
$\tilde{s}^{\prime}_{i}=\tilde{s}_{i}+\alpha\frac{n}{\|n\|},$ (4)
where $\alpha$ is a scaling factor. While it applies uniform addition across vectors, its effect varies in the cosine similarity computation between $\tilde{s}^{\prime}_{i}$ and ${s}_{i}$. The adjustment automatically emphasizes loss for safe vectors with lower correlation to $n$ and reduces it for those with higher correlation. Figure 4 illustrates this behavior: a safe vector with low initial correlation to $n$ incurs a larger loss, while one with higher correlation yields a smaller loss. Integrating this adjustment with the loss that minimizes the cosine similarity between $s_{i}$ and $\tilde{s}_{i}$, we propose the safe loss as:
@@
$\mathcal{L}_{s}\!=\!\frac{1}{B}\!\sum_{i=1}^{B}\!\left[\!\!\biggl(\!1\!-\!\frac{\tilde{s}_{i}\cdot{s}_{i}}{\|\tilde{s}_{i}\|\!\|{s}_{i}\|}\!\biggr)\!+\!\biggl(\!1\!-\!\frac{\tilde{s}^{\prime}_{i}\cdot{s}_{i}}{\|\tilde{s}^{\prime}_{i}\|\!\|{s}_{i}\|}\!\biggr)\!\!\right]\!,\!$ (5)
where each of $s_{i}$, $\tilde{s}_{i}$, and $\tilde{s}^{\prime}_{i}$ represents an $i$-th embedding vector in a mini-batch. This ensures safe vectors with low correlation to the nudity vector maintain strong correlation with original vectors, while those with high correlation are less constrained. Thus, distinctly safe embeddings retain their semantics, while ambiguous ones are moderately adjusted with unsafe embeddings through the unsafe loss. This adaptive behavior allows flexible embedding space distortion while preserving clearly safe embeddings.
3.2.3 Nudity Embedding Neutralization
Furthermore, there might still be attempts to exploit nudity embedded in text encoders. For instance, Ring-A-Bell extracts the nudity vector and uses genetic algorithms to find prompts whose embeddings are similar to the combination of safe embeddings and the extracted concept. To prevent such extraction-based attacks, we propose the nudity neutralization loss, which aims to neutralize the semantic meaning of the nudity vector itself. We achieve this by aligning the “nudity” vector with the neutral vector (i.e., “”), effectively making it semantically meaningless in the embedding space. Nudity neutralization loss is represented as:
$\mathcal{L}_{n}=1-\frac{\tilde{n}\cdot{e}_{0}}{\|\tilde{n}\|\|{e}_{0}\|},$ (6)
@@ 3.2.3 Nudity Embedding Neutralization
where $\tilde{n}$ and ${e}_{0}$ denote the current “nudity” vector and the neutral vector, respectively. This alignment ensures that even if adversaries attempt to extract the “nudity”, they will only obtain a semantically neutral embedding that cannot be effectively used for attacks. Thus, while the unsafe loss provides robustness against adversarial attacks, nudity neutralization loss complements this by eliminating the possibility of direct concept exploitation.
@@
where $\lambda$ controls the balance between unsafe, safe, and nudity neutralization losses to distort the unsafe embedding space while preserving the safe embeddings. (Analysis of $\lambda$ can be found in Appendices D.3 and D.4.) Note that the nudity neutralization loss operates on the current “nudity” vector from the training text encoder, while the unsafe loss transforms unsafe vectors into safe regions defined using a pre-computed “nudity” vector subtracted from selected safe vectors. Similarly, the loss adjustment in the safe loss also uses the pre-computed “nudity” vector for similarity calculation. These three losses are therefore complementary and do not conflict with each other. We present the overall DES training process in Algorithm 2.
@@
0: Original text encoder $\mathcal{E}_{\phi_{o}}$, paired set $\mathcal{D}$, nudity vector $e_{n}$, neutral prompt $p_{0}$, scale factor $\alpha$, hyperparameter $\lambda$, mini-batch size $B$, iteration $T$
[…]
@@
5:  $(t,p_{u},p_{s})\leftarrow$ Read one mini-batch from $\mathcal{D}$
6:  $s\leftarrowRead$ one mini-batch from $\mathcal{S}$
7:  $\tilde{u},\tilde{s}\leftarrow\mathcal{E}_{\phi}(p_{u}),\mathcal{E}_{\phi}(p_{s})$
8:  $\tilde{s}^{\prime}_{i}$ is computed by Eq. (4) for $i=1,\ldots,B$
9:  $\tilde{n}\leftarrow\mathcal{E}_{\phi}(p_{n})$ // Extract current nudity vector
10:  Total loss $\mathcal{L}_{t}$ is computed by Eq. (7) using $t,\tilde{u},s,\tilde{s},\tilde{s}^{\prime},\tilde{n},e_{0}$
@@ 4.1 Experimental Settings
Baseline Models. Our experiments are conducted on SDv1.4 and v1.5 [stablediffusion], widely adopted open-source models [guardt2i, advunlearn], and FLUX.1 [flux], a recently introduced popular model, for the T2I tasks. Additionally, we use SD-inpainting, which takes a mask image as an additional input, for the I2I tasks. We set $\lambda=0.3$ and $\alpha=200$ to train the text encoders of SDv1.5 and FLUX.1. Since FLUX.1 uses multiple text encoders, we train each encoder independently using the same settings. Further implementation details, evaluation on additional models, ablation studies are provided in Appendices A, B.4, D.1, and D.2.
Threat Models. We evaluate DES and other approaches under three threat scenarios: explicit prompts, black-box adversarial prompts, and white-box adaptive attacks. For explicit prompts, we use the I2P dataset, which may be created intentionally or unintentionally by users without model access. For black-box attacks, where attackers lack model access but rely on prompt engineering or transferability, we use prompts like Sneaky, MMA, Ring-A-Bell, and P4D. For white-box scenarios, where attackers have full model access and use optimization-based methods, we evaluate against UDA [unlearndiffatk], Ring-A-Bell, MMA, and CCE [cce]. All evaluations use publicly available unsafe prompts.
Training Datasets. For Sections 4.2 and 4.3, we use 6,911 safe–unsafe prompt pairs from the sexual category of CoPro dataset [latentguard] to train the text encoder. For Section 4.4, we additionally use 8,931 prompt pairs from the violence and illegal categories to cover NSFW categories such as violence, illegal, hate, and others. We also generate 1,600 prompts related to Van Gogh for experiments in Section 4.4.
Comparison Models. We compare DES against other approaches, such as OpenAI moderation [openai_moderation], Microsoft Azure [azure], Latent Guard, GuardT2I, SAFREE, SLD [sld], SPM [spm], UCE, ESD, GLoCE [gloce], and SalUn. We also include text encoder-based approaches such as Safe-CLIP [safeclip] and AdvUnlearn for direct comparison with DES.
Metrics. We evaluate our method using three main metrics. ASRs in Section 4.2 and 4.3 are measured using NudeNet [nudenet], a nudity detector, while ASRs in Section 4.4 are measured using Q16 [q16] to cover a broader range of unsafe content. Image generation quality is assessed using FID, and text–image alignment is evaluated using CLIP score [clipscore], computed on 10k samples from COCO 30k dataset [coco].
Table 1: Quantitative comparison of defense methods against I2P prompts in T2I using SDv1.5. NudeNet is utilized to detect nudity, with female and male body parts denoted as (F) and (M), respectively. The best and second-best scores are highlighted in red and blue, respectively.
[…]
SDv1.5 196 30 47 34 62 76 183 223 851 16.57 26.46
SPM 153 25 37 34 49 60 143 203 704 16.65 26.46
SLD-strong 65 7 54 30 47 52 117 139 511 31.38 24.61
Safe-CLIP 89 8 28 5 24 35 84 131 404 17.49 25.73
SAFREE 26 1 37 17 18 41 57 66 263 27.09 25.82
UCE 31 1 18 14 15 21 60 56 216 16.99 26.16
ESD 16 0 5 3 4 17 23 37 105 17.75 25.30
GLoCE 22 9 3 1 6 14 27 23 105 21.21 25.70
SalUn 0 0 0 2 0 14 1 4 21 21.14 24.78
AdvUnlearn 1 0 1 0 2 5 5 13 27 18.94 23.82
DES (ours) 1 0 0 0 0 7 3 5 16 15.44 25.52
@@ 4.2 Experimental Results on T2I
4.2.1 Explicit Sexual Content Mitigation
To evaluate mitigation performance against explicit sexual content, we measure the number of nude body parts in generated images using NudeNet, as shown in Table 1. Existing methods frequently generate unsafe content, such as breasts, genitalia, and buttocks. SPM and SLD show particularly poor performance, likely because they preserve model parameters, allowing nudity to remain. SAFREE and GLoCE perform better but still suffer from retained parameters, leaking content such as female breasts. In contrast, most nudity detected in DES consists of relatively safe body parts, such as feet, belly, and armpits, while generating only one instance of female breasts. Even when including these, DES shows SOTA performance with only 16 total detections.
Although AdvUnlearn and SalUn also reduce explicit content, both face notable limitations in benign
Table 2: Quantitative comparison of defense methods against adversarial prompts in T2I using SDv1.5 and FLUX.1. Models marked with are evaluated using filtering accuracy instead of NudeNet. The best and second-best scores are highlighted in red and blue, respectively.